OpenStudy (watchmath):

Compute the exact value of \[\int_0^{\pi/2}\frac{\sin x}{9+16\sin(2x)}\] (wolfram alpha will give you an approximation to this integral :D)

OpenStudy (anonymous):

holy pellet. no way im trying this :P

OpenStudy (anonymous):

Before I even attempt it, can it be done? (I mean, its not a something like e^(-x^2)

OpenStudy (watchmath):

Sure it is solvable :D

OpenStudy (anonymous):

Hmmm...well let me attempt it.

OpenStudy (anonymous):

If you know how to do that. Please show me. :P

OpenStudy (watchmath):

Use substitution $u=\pi/2 -x$ then use this : http://openstudy.com/users/watchmath#/users/watchmath/updates/4dd3bbd699508b0bc2d6a5a4 It is tough!

OpenStudy (watchmath):

I mean \(u=\frac{\pi}{2}-x\)

OpenStudy (anonymous):

watchmath!!! Check my response to that link you posted please D:

OpenStudy (watchmath):

:)

OpenStudy (anonymous):

You suck D;